OpenStudy (broskishelleh):

h For a right triangle with legs measuring seventeen 17 and 9, a student submitted the following explanation of how to find the exact length of the hypotenuse. Explain the student's error.

OpenStudy (broskishelleh):

only 18 on the picture

OpenStudy (broskishelleh):

I fanned you ganibl

OpenStudy (shamil98):

he substituted the one of the legs as c, which is supposed to be the hypotenuse.

OpenStudy (shamil98):

\[\huge a^2 + b^2 = c^2\] He should have put: \[\huge 17^2 + 9 ^2 = c^2\]

OpenStudy (shamil98):

Basically, he incorrectly placed the value of 17^2 as c^2, c = hypotenuse 17 and 9 are the legs (which are a and b)

OpenStudy (broskishelleh):

Thanks:)